OpenStudy (anonymous):

Fan Testimonial and Medal! What is the slope of the line that passes through the points (1, 7) and (–4, –8)? What is the slope of the line that passes through the points (–2, 15) and (–8, –5)?

OpenStudy (anonymous):

Only the second line!

OpenStudy (anonymous):

If you can solve 1, then you can solve second one also..

Vocaloid (vocaloid):

point slope formula: (y2-y1) = m(x1-x2) solve for m by using the points (x1,y1) and (x2,y2) by plugging in the values for x1,x2,y1, and y2

OpenStudy (anonymous):

i didn't line one is a different question lol

OpenStudy (anonymous):

I know, but they are one and the same concept based question. If you can find the slope of first, then by using same formula, you can find second one too. :P

OpenStudy (anonymous):

i didn't the first line is the wrong problem (the next question)

OpenStudy (anonymous):

Slope is given by: \[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

so 3/17?

OpenStudy (anonymous):

How, show me first..

OpenStudy (anonymous):

y2-y1 =3 ------- x2-x1= 17

OpenStudy (anonymous):

Okay, what is y2 value and what is y1 value there?

OpenStudy (anonymous):

-5 minus -8

OpenStudy (anonymous):